OpenStudy (anonymous):

you have 4/6 cup of brown sugar in your cupboard .The recipe for a dessert calls for 2/5 cup of brown sugar. How much brown sugar would you have left if you made the dessert help plz >??

OpenStudy (anonymous):

4/6-2/5=2/3-2/5=10/15-6/15=4/15